# The Dielectric Response of Plasmas with Arbitrary Gyrotropic Velocity Distributions

### v1

@@ Appendix B Eigenfunctions and Heating Rates
We can then determine the perturbed velocity fluctuations for each component $\delta\bm{U}_{s}$ by recognizing that the total current density (including any parallel flow) due is $\bm{j}_{s}=q_{s}(n_{0s}\,\delta\bm{U}_{s}+\delta n_{s}U_{s}\mbox{$\hat{\mathbf{z}}$})$. Using the susceptibility tensor to calculate $\bm{j}_{s}$ through Eqn. 1 yields
@@ Appendix C The Limits of Analytic Continuation
Given the novel GLLS representation for the analytic continuation, it is important to quantify how far into imaginary parallel momentum space the representation can be accurately and reliably extended. Towards this end, we consider the continuations of three VDFs: a single Maxwellian, the best fit core-and-beam bi-Maxwellian $f_{C\&B}(p_{\perp}=0)$ used in § III.1, and the observed VDF $f_{Obs.}(p_{\perp}=0)$ also described in § III.1. We evaluate the amplitude of the complex valued representation of the VDF $|f_{p}[\mathbb{R}e(p_{\parallel}),\mathbb{I}m(p_{\parallel});p_{\perp}=0]$ for all three models. For the GLLS representation, we apply orders 5, 10, 30, and 50. These comparisons are shown in Fig. 8, and the coefficients for the GLLS fits are shown in Fig. 9.
@@
This increase in $|f_{p}|$ impacts the resolution of strongly damped waves, as the associated values of $|\mathbb{I}m(p_{\parallel})|$ for such modes are large enough for the numerical evaluation of the analytic continuation of $f_{p}$ to be inaccurate. However, given the superior behavior of the GLLS method for weakly damped solutions, we find using the Chebyshev polynomials rather than a low-order Maxwellian fit is the optimal choice for non-Maxwellian VDFs.
Figure 8: The amplitude of $|f_{p}|$ evaluated at complex parallel momentum for $p_{\perp}=0$ using one or two bi-Maxwellians, as well as for several orders of GLLS (5 through 50; right). The input VDFs are: (left) a single Maxwellian, (center) the core-and-beam bi-Maxwellian from § III.1, and (right) the observed distribution from § III.1.

### v2

@@ Appendix B Eigenfunctions and Heating Rates
[…]
@@ Appendix C The Limits of Analytic Continuation
Given the novel GLLS representation for the analytic continuation, it is important to quantify how far into imaginary parallel momentum space the representation can be accurately and reliably extended. Towards this end, we consider the continuations of three VDFs: a single Maxwellian, the best fit core-and-beam bi-Maxwellian $f_{C\&B}(p_{\perp}=0)$ used in § IV.1, and the observed VDF $f_{Obs.}(p_{\perp}=0)$ also described in § IV.1. We evaluate the amplitude of the complex valued representation of the VDF $|f_{p}[\mathbb{R}e(p_{\parallel}),\mathbb{I}m(p_{\parallel});p_{\perp}=0]$ for all three models. For the GLLS representation, we apply orders 5, 10, 30, and 50. These comparisons are shown in Fig. 9, and the coefficients for the GLLS fits are shown in Fig. 10.
@@
This increase in $|f_{p}|$ impacts the resolution of strongly damped waves, as the associated values of $|\mathbb{I}m(p_{\parallel})|$ for such modes are large enough for the numerical evaluation of the analytic continuation of $f_{p}$ to be inaccurate. As seen in the right-hand panels of Fig. 9, there is a larger increase in $|f_{p}|$ with higher order representations for more strongly damped solutions. For this reason, users should select the lowest-order representation sufficient to capture the VDF structure. This approach not only optimizes the numerical efficiency of this scheme, but it also avoids using unnecessarily high orders which may introduce numerical artifacts for strongly damped solutions. Overall, given the superior behavior of the GLLS method for weakly damped solutions, we find using the Chebyshev polynomials rather than a low-order Maxwellian fit is the optimal choice for non-Maxwellian VDFs.
Figure 9: The amplitude of $|f_{p}|$ evaluated at complex parallel momentum for $p_{\perp}=0$ using one or two bi-Maxwellians, as well as for several orders of GLLS (5 through 50; right). The input VDFs are: (left) a single Maxwellian, (center) the core-and-beam bi-Maxwellian from § IV.1, and (right) the observed distribution from § IV.1.